INFRARED SPECTROSCOPY OF PROTONATED METHANOL-WATER CLUSTERS -EFFECTS OF HET-EROMOLECULES IN HYDROGEN BOND NETWORK-

\underline{KEN-ICHIRO SUHARA}, ASUKA FUJII and NAOHIKO MIKAMI, Department of Chemistry, Graduated School of Science, Tohoku University, Sendai 982-8578, Japan.

IR spectroscopy of H^{+}(MeOH)_{m}(H_{2}O)_{n}(m = 2 - 13, n = 1 and 2) was carried out in the OH stretch region to characterize the development of the hydrogen bond network with the cluster size. For n = 1, the band intensity of the free OH stretching mode decreased with the cluster size, and the band finally disappeared at m = 7. This result is reasonably interpreted in terms of the formation of the tricyclic structure of the hydrogen-bond network in the clusters.